An Engineer Invests $2000 In An Individual Retirement Account (IRA) Each Year During Her 40 Year Career. Assume The Investment Earns An 8% rate of return. what annual income can be drawn from the investment for the 30 year following her retirement? A) 46000 B)2670 C) 17300 D)4500 E) none of the above
Answer #1

Yearly investment = 2000

Time period = 40 years

Rate of return = 8%

What annual income can be drawn from the investment for the 30 year following her retirement?

Step 1

Calculate Future Value of the investments

FV = 2000 (F/A, 8%, 40)

FV = 2000 (259.05651)

FV = 518,113

FV at 40th year = 518,113

Let this amount is the Present value

Now calculate annual income can be drawn from the investment for the 30 year

A = P (A/P, 8%, 30)

A = 518,113 (A/P, 8%, 30)

A = 518,113 (0.08883) = 46,023.9 or 40,024

The answer nearer to this value given in the option is 46,000

Answer - A) $46,000